# Impedance vs. Power Side-channel Vulnerabilities: A Comparative Study

### v1

@@ Abstract
In recent times, impedance side-channel analysis has emerged as a potent strategy for adversaries seeking to extract sensitive information from computing systems. It leverages variations in the intrinsic impedance of a chip’s internal structure across different logic states. In this study, we conduct a comparative analysis between the newly explored impedance side channel and the well-established power side channel. Through experimental evaluation, we investigate the efficacy of these two side channels in extracting the cryptographic key from the Advanced Encryption Standard (AES) and analyze their performance. Our results indicate that impedance analysis demonstrates a higher potential for cryptographic key extraction compared to power side-channel analysis. Moreover, we identify scenarios where power side-channel analysis does not yield satisfactory results, whereas impedance analysis proves to be more robust and effective. This work not only underscores the significance of impedance side-channel analysis in enhancing cryptographic security but also emphasizes the necessity for a deeper understanding of its mechanisms and implications.
@@
Impedance side-channel, side-channel analysis, impedance leakage, key extractions, side-channel attack
@@ I Introduction
In cybersecurity, side-channel attacks have long been recognized as potent threats, exploiting unintended information leakage to compromise the security of cryptographic systems and electronic devices. Traditional side channels, such as power, electromagnetic (EM) emissions, and timing variations, have received significant attention in academia and industry due to their ability to leak sensitive information from cryptographic devices and systems [1, 2, 3, 2, 4, 5, 6]. For instance, power side-channel attacks leverage fluctuations in the power consumption of cryptographic devices across different operations, such as encryption and decryption processes. Similarly, EM emissions arise from the electrical operations within cryptographic devices, leading to unintentional radiations that attackers can intercept and analyze [4, 5]. Usually, the vulnerabilities in the hardware implementation or the physical characteristics of the encryption systems facilitate these unintended physical side channels, which are then exploited to extract confidential information from cryptographically secure algorithms [1, 2, 3, 2, 4, 5, 6].
@@
Recently, impedance analysis has found applications across various domains, including radio frequency (RF) related fields [7], counterfeit chip detection [8] and tempering detection [9]. It has also been discovered that impedance can inadvertently leak sensitive information related to the computations and data being processed within a system. Our research demonstrates the potential of the runtime impedance of a device as a side-channel, unveiling software instructions through distinct impedance profiles generated by each operation [10, 11]. Subsequent studies have exploited impedance as a side channel to extract the AES key from the register of an FPGA [12]. These approaches to leveraging impedance marks a significant shift in the perception of impedance within the cybersecurity domain. Impedance analysis, while it remains an important tool for enhancing hardware security, it also presents a novel vulnerability that adversaries can exploit to their advantage.
Although studies demonstrate the extraction of the cryptographic key, such as the AES key, using power and impedance side channels, there remains a gap in the literature regarding the comparative evaluation of these two side channels. This study endeavors to bridge this gap by providing a detailed comparative analysis of the performances of impedance and power side channels in the extraction of the AES key. The primary contributions of this paper are,
@@
  • Exploring the source of impedance side-channel leakage and the extraction of AES keys using it. This includes scenarios with and without countermeasures.
@@
  • Performing a comparative analysis of the performances of impedance and power side channels.
@@
The structure of this paper is organized as follows: Section II discusses the background of power and impedance side channels, including a discussion on correlation analysis and the existing research on the extraction of encryption keys using physical side channels. Section III presents the proposed method to compare the power and impedance side channels. Section IV details the experimental setup. We present and compare the findings of AES key extraction using power and impedance side channels in Section V. Section VI concludes the work.
@@ II Background and Related Work
This section covers the basics of the cryptographic algorithm targeted in this study, explores power and impedance side channels, and outlines the attack methodology used. Furthermore, it reviews recent literature on existing power and impedance side channel-based work.
@@ II-A Advanced Encryption Standard
This study focuses on the advanced encryption standard (AES), a commonly used symmetric block cipher algorithm that encrypts and decrypts data using the same key. Figure. 1 presents the overview of AES encryption steps for $r$ rounds. Each round involves a key-dependent substitution step using the Rijndael S-BOX [13, 14], a step that shifts rows, a step that mixes columns, and a step that adds a round key [15]. The Rijndael S-BOX is a lookup table and is critical for the security of the algorithm. It changes an 8-bit input to an 8-bit output, making the encryption stronger by ensuring that a small change in the input significantly affects the output. This study considers the AES-128 algorithm to compare how well impedance and power side channels can be used to analyze it.
@@ II-B Power Side Channel
The power side channel focuses on the power consumption of the target devices and is a well-studied area [1]. It originates from the fact that, in digital circuits, the execution of operations depends on the switching activities of the transistors. These switching activities lead to voltage and current fluctuations across digital components, resulting in power consumption variations. These fluctuations can be captured by measuring the voltage changes across the power supply pins of the target component using an oscilloscope. The resultant measurements can be subsequently used in further analysis and processing to extract information related to the operation being executed. This poses a security and privacy issue as it can be exploited to obtain sensitive credentials.
This work focuses on using power side channels to extract the AES key. We explore a scenario within an embedded system that requires the implementation of AES. We study the internal memory modules to be read and written with the intermediate values of the computations performed in the encryption process. The dynamic power consumption of these memory modules, corresponding to read and write operations, is intrinsically linked to the changes in the logic states of the transistors. During the write operation, we focus on the state transitions of the transistors, either transitioning from ON to OFF or vice versa. By measuring and analyzing these power consumption patterns, we aim to recover the AES encryption key.
@@ II-C Impedance Side Channel
In contrast to the power side channel, the impedance side channel utilizes changes in the intrinsic impedance variation of the system based on different logic states. The logic state of the device represents various transistor configurations, resulting in impedance variations [16, 10]. The impedance variation due to stored binary values in memory cells exhibits distinct differences. This variation is evident when writing specific data to the memory across different clock cycles, where the number of transistors in the ON and OFF states varies and remains constant until the next clock signal. By exploiting this phenomenon, our focus shifts to the impedance variations resulting from the storage of intermediate values of AES encryption in the memory chip. We explore the use of impedance measurements to extract the encryption key, leveraging these variations.
To measure the impedance, we use a Vector Network Analyzer (VNA). VNA is widely used to characterize multi-port networks. The VNA operates by injecting a known test signal and analyzing the reflected and transmitted signals from the device under test (DUT). This process allows the VNA to calculate the DUT’s scattering parameters, commonly referred to as S-parameters, which are widely used in the fields of radio frequency (RF) and microwave systems. A key focus of our analysis is the input port reflection coefficient, $S_{11}$, from which we derive the supply side impedance of the DUT, $Z_{i}$, using 1. Here, $Z_{ref}$ represents the reference impedance.
@@ II-D Correlation Analysis
Recent works use different types of analytical methods to signals in physical side channels. One popular analysis method is correlation-based analysis. This analysis is considered powerful in unmasking the relationships of the secret information hiding inside the noisy measurements using the Pearson correlation coefficient as a statistical test [17]. Furthermore, instead of focusing on only a single measurement point, it adopts a multivariate approach by analyzing multiple points within the measurement trace [1].
@@
In cryptography, correlation-based analysis starts with the collection of leakage responses generated by known plaintext operations. Each measurement is synchronized with the cryptographic encryption operation performed in the device over the plaintext. The next step involves selecting a leakage model that represents hypothesized side-channel leakages. This leakage model is then correlated with the actual observed leakage to find the leaked information. To facilitate this analysis, let $M$ represent the total number of measurements, $S$ the total number of sampling points per measurement, $x_{s,m}$ the actual leakage at the $s$’th sample point of the $m$’th measurement, and $h_{k,m}$ the hypothesized leakage for a guessed key $k$ at the $m$’th measurement. The relationship between the actual and hypothesized leakages is quantified by the Pearson correlation coefficient $\rho{(x_{s},h_{k})}$, calculated as presented in 2.
@@ II-E Related Work
Power side-channel analysis on both the hardware and software implementations of cryptographic algorithms has been extensively researched over the past few decades. Authors in [18] present one of the earliest power side-channel attacks to extract the AES key from smart cards. In [19], the authors discuss a detailed explanation of step-wise differential power analysis (DPA) attacks over AES implementation. In [20], three attack scenarios are presented: single-bit DPA, multi-bit DPA, and a correlation analysis. The authors implement AES in two platforms. The study concludes by proposing an enhanced DPA approach that strategically organizes plaintext inputs to maximize the leakage. Additionally, [21, 22] present AES key extraction attacks using power analysis on the ATmega328P microcontroller. Further, [23] presents a power analysis attack specifically targeted against ASIC implementations of the AES. The authors in [24, 25] present their findings on power side-channel attacks targeting the advanced modes of AES.
Impedance is a recent addition to the field of side-channel analysis [26, 11, 27, 16, 10, 26]. Before its application to cryptographic key recovery, numerous studies focused on using impedance for hardware security. Authors in [28] introduce a detection technique using board level impedance to detect physical modifications of the board. The study in [26] explores a method for non-destructively detecting hardware Trojans through internal impedance variations. The authors in [29], study a clustering method to classify ICs into Trojan-free and Trojan-infected groups. In addition to the usage of impedance in physical modification detection, it is recently attracting the focus of device switching activities. Authors in [16] use impedance measurements to disassemble software instruction types for anomaly monitoring and software integrity verification. The study of impedance changes induced by the switching activities is used to detect and identify the firmware in [11]. The authors in [30] use the impedance of digital circuits to introduce a novel RFID tag design. These works present the run-time impedance variations of the devices as the side-channel.
@@ III Power vs Impedance: Key Extraction
This section discusses the attack methodology performed against the AES implementation utilizing power and impedance side channels. We present the threat model, the leakage model used in the attack, and the metrics to evaluate the power and impedance side channels.
@@ III-A Source of Impedance Side-channel
Impedance side-channel analysis introduces a novel approach to probing the intricate and often overlooked behaviors of electronic circuits, particularly in memory chips. This investigative technique methodically explores memory operations, revealing how variations in electrical impedance could potentially disclose information about the data being processed and stored. At the core of this analysis lies the understanding that various physical and architectural factors within a circuit influence its overall impedance characteristics [16, 26]. These factors include parasitic elements inherent to physical implementations, interconnection structures, transistor property variations across the silicon substrate, and the distinct semiconductor material properties of the transistor configurations.
Within memory chips, such as Static Random Access Memory (SRAM) and Dynamic Random Access Memory (DRAM), the control circuit manages memory access and data read/write operations. The memory array, organized in rows and columns, stores data at specific address locations. Three primary components define the impedance profile of memory: the architecture and layout of the memory array, the cell type, and the related data and control logic circuitry. Manufacturers tailor the arrangement of memory cells in rows and columns, as well as the cell type (e.g., four-transistor (4T) or six-transistor (6T) cell structures in SRAM), based on factors such as performance requirements, power consumption, and area constraints, thereby influencing the impedance characteristics.
However, in accordance with our study, we focus on the impedance changes corresponding to the memory content. The process of writing data to or reading from memory cells initiates a series of events that create unique impedance profiles for the stored bit sequences within the memory array. Specifically, the activation of transistors, fabrication process variations, and the specific data retention logic employed contribute to these distinctive impedance profiles. Furthermore, the impedance profiles are influenced by factors such as the data stored inside the memory cells, the different locations of the cells in terms of rows and columns within the array, and the varying lengths of connections for each cell. Thus, the impedance measurements should reflect these variations and can be vulnerable to data leakage.
@@ III-B1 Target Intermediate
The key extraction of the AES encryption using physical side-channel analysis usually targets the first or final round of the algorithm. The targeted round is used to model the leakage signal. In our work, we target the first round of the AES algorithm. The AES encryption begins with performing the bitwise XOR operation on the plaintext and the secret key, and the resulting product is used in the substitution step. The output value of the substitution step is stored in memory, and it is our targeted computational intermediate value. The steps of AES encryption are illustrated in 1.
We use the plaintext in our hypothetical leakage model. With the known plaintext $p_{i}$, and the corresponding round subkey $k_{j}$, the targeted computational intermediate $m_{i}$ can be represented as in 3. This intermediate value depends on both the input plaintext and the secret cryptographic key. Thus, the corresponding side-channel leakages are correlated with the secret key information. The attackers can exploit the statistical properties of the leakage traces to extract the cryptographic key.

### v3

@@ Abstract
Physical side channels emerge from the relation between internal computation or data with observable physical parameters of a chip. Previous works mostly focus on properties related to current consumption such as power consumption. The fundamental property behind current consumption occur from the impedance of the chip. Contemporary works have stared using chip impedance as a physical side channel in extracting sensitive information from computing systems. It leverages variations in intrinsic impedance of a chip across different logic states. However, there has been a lack of comparative studies. In this study, we conduct a comparative analysis of the impedance side channel, which has been limitedly explored, and the well-established power side channel. Through experimental evaluation, we investigate the efficacy of these side channels in extracting stored advanced encryption standard (AES) cryptographic key on a memory and analyze their performance. Our findings indicate that impedance analysis demonstrates a higher potential for cryptographic key extraction compared to power side-channel analysis (SCA). Moreover, we identify scenarios where power SCA does not yield satisfactory results, whereas impedance analysis proves to be more robust and effective. This work not only underscores the significance of impedance SCA in enhancing cryptographic security but also emphasizes the necessity for a deeper understanding of its mechanisms and implications.
@@
Impedance side-channel, side channel analysis, impedance leakage, key extractions, side channel attack
@@ I Introduction
In cybersecurity, side-channel attacks have long been recognized as potent threats, exploiting unintended information leakage to compromise the security of cryptographic systems and electronic devices. Traditional side channels, such as power, electromagnetic (EM) emissions, and timing variations, have received significant attention in academia and industry due to their ability to leak sensitive information from cryptographic devices and systems [1, 2, 3, 4, 5, 6]. For instance, power side-channel attacks leverage fluctuations in the power consumption of cryptographic devices across different operations, such as encryption and decryption processes. Similarly, EM emissions arise from the electrical operations within cryptographic devices, leading to unintentional radiations that attackers can intercept and analyze [4, 5]. Usually, the vulnerabilities in the hardware implementation or the physical characteristics of the encryption systems facilitate these unintended physical side channels, which are then exploited to extract confidential information from cryptographically secure algorithms [1, 2, 3, 4, 5, 6].
@@
Recently, impedance analysis has found applications across various domains, including radio frequency (RF) related fields, counterfeit chip detection and tempering detection [7, 8, 9]. It has also been discovered that impedance can inadvertently leak sensitive information related to the computations and data being processed within a system. Our research demonstrates the potential of the runtime impedance of a device as a side-channel, unveiling software instructions through distinct impedance profiles generated by each operation [10, 11]. Subsequent studies have exploited impedance as a side channel to extract the AES key from the register of an FPGA [12]. These approaches to leveraging impedance marks a significant shift in the perception of impedance within the cybersecurity domain. Impedance analysis, while it remains an important tool for enhancing hardware security, it also presents a novel vulnerability that adversaries can exploit to their advantage.
Despite the extensive focus on current and voltage as the primary sources of side-channel leakage, impedance, the parameter that connects these two, offers a distinct and underexplored perspective for SCA. Unlike power-based techniques, which primarily measure current fluctuations, impedance captures the dynamic interaction between current and voltage, providing a distinct and potentially more insightful view into data-dependent variations within a system. Although impedance has indirectly contributed to static power analysis attacks, its direct significance has often been overlooked [13, 14, 15]. This is primarily due to the conventional assumption that impedance remains constant, determined by the chip’s physical structure, dimensions, and fabrication materials. However, the influence of operational state changes on impedance, and their potential contribution to side-channel leakage, has not been comprehensively investigated. This paper addresses this gap by investigating the differences between impedance and power side channels, with a focus on demonstrating the effectiveness of impedance in cryptographic key extraction. Through a comprehensive analysis, this study aims to highlight the potential of impedance as a side-channel vector, offering a new perspective on hardware vulnerability assessment. The primary contributions of this paper are,
  • Investigating the sensitivities of impedance and power in SCA. Although these side channels are interconnected, this research clarifies how they can exhibit differing sensitivities.
@@
  • Exploring impedance side-channel leakage sources and AES key extraction using this method, including scenarios with and without the presence of background noise.
@@
  • Performing a comparative analysis of the performance of impedance and power side channels.
@@
The structure of this paper is organized as follows: Section II discusses the background of power and impedance side channels, including a discussion on correlation analysis and the existing research on the extraction of encryption keys using physical side channels. Section III presents the proposed method to compare the power and impedance side channels. Section IV details the experimental setup. We present and compare the findings of AES key extraction using power and impedance side channels in Section V. Section VII concludes the work.
@@ II Background and Related Work
This section provides an overview of the cryptographic algorithm targeted in this study and the concepts of power and impedance side channels. It then outlines the attack methodology used in our analysis. Additionally, recent literature on power and impedance side-channel research is reviewed to contextualize the existing work and highlight the contributions of this study.
@@ II-A Advanced Encryption Standard
This study focuses on the AES, a widely adopted symmetric block cipher that encrypts and decrypts data with the same key. Fig. 1 presents the AES encryption process, which consists of $r$ rounds. Each round includes key-dependent operations: substitution using the Rijndael S-BOX [16, 17], row shifting, column mixing, and round key addition [18]. The Rijndael S-BOX, implemented as a lookup table, plays a critical role in security by transforming an 8-bit input into an 8-bit output, ensuring that small changes in input lead to significant alterations in the output. In this work, the AES-128 algorithm is used to compare the effectiveness of power and impedance side channels in cryptographic analysis.
@@ II-B Power Side Channel
The power side channel exploits the power consumption patterns of target devices and is a well-established area of research [1]. In digital circuits, power consumption is directly related to the switching activities of transistors during computation. These switching events cause fluctuations in current, leading to measurable variations in power consumption. By capturing voltage and current changes at the power supply pins of the target device using an oscilloscope, these variations can be recorded for further analysis. Such measurements provide indirect information about the executed operations, potentially exposing sensitive data like cryptographic keys, which poses significant security and privacy risks.
This work focuses on exploiting power side channels to extract the AES encryption key. We target an embedded system running AES. We investigate the dynamic power consumption of internal components, such as memory modules, where intermediate values are temporarily stored during the encryption process. The power consumption during memory read and write operations is inherently linked to the switching activities of transistors as they transition between ON and OFF states. However, rather than directly measuring these transitions, power side channels capture the cumulative impact of switching activities on the overall power consumption. By analyzing these patterns, particularly during write operations when state transitions occur, we aim to recover the AES encryption key, highlighting the vulnerability of embedded systems to power-based side-channel attacks.
@@ II-C Impedance Side Channel
In contrast to power side channel, impedance side channel exploits variations in a system’s intrinsic impedance, which fluctuates with different logic states. These states correspond to various configurations of transistors, leading to measurable impedance changes [19, 10]. Such variations are particularly pronounced in memory cells, where distinct impedance differences correspond to the binary data stored. This effect is evident across clock cycles in edge-triggered synchronous systems, as the number of transistors in ON and OFF states remains constant until the next clock signal. By exploiting these impedance variations, we focus on analyzing the storage of intermediate values during AES encryption in memory chips. This approach enables the extraction of cryptographic keys through impedance measurements.
@@
To measure the impedance, we use a vector network analyzer (VNA). VNA is widely used to characterize multi-port networks. The VNA operates by injecting a known test signal into the device under test (DUT) and analyzing the reflected and transmitted signals. Through this process, the VNA calculates the DUT’s scattering parameters (S-parameters), which are important to RF and microwave system analysis. Our analysis focuses on the input port reflection coefficient, $S_{11}$, from which we derive the supply side impedance of the DUT, $Z_{i}$, using Eq. 1, where $Z_{ref}$ denotes the reference impedance. These impedance measurements form the basis of our work, allowing us to evaluate the feasibility of using impedance side channels for cryptographic key extraction.
@@ II-D Correlation Analysis
Recent works use various analytical methods to study signals in physical side channels. One popular analysis method is correlation-based analysis. This analysis is considered powerful in unmasking the hidden relationships between secret information and noisy measurements using the Pearson correlation coefficient as a statistical test [20]. Additionally, rather than relying on a single measurement point, correlation-based analysis uses a multivariate approach by analyzing multiple points within the measurement trace, thereby improving the robustness and accuracy of the side-channel attack [1].
@@
Correlation-based analysis begins with the collection of leakage responses generated during known plaintext operations. Each measurement is synchronized with the cryptographic encryption operation performed in the device over the plaintext. The next step involves selecting a leakage model that represents hypothesized side-channel leakages. This leakage model is then correlated with the actual observed leakage to find the leaked information. To facilitate this analysis, let $M$ represent the total number of measurements, $S$ the total number of sampling points per measurement, $x_{s,m}$ the actual leakage at the $s$’th sample point of the $m$’th measurement, and $h_{k,m}$ the hypothesized leakage for a guessed key $k$ at the $m$’th measurement. The relationship between the actual and hypothesized leakages is quantified by the Pearson correlation coefficient $\rho{(x_{s},h_{k})}$, calculated as presented in Eq. 2.
@@ II-E Related Work
Power SCA on both the hardware and software implementations of cryptographic algorithms has been extensively researched over the past few decades. Authors in [21] present one of the earliest power side-channel attacks to extract the AES key from smart cards. In [22], the authors discuss a detailed explanation of step-wise differential power analysis (DPA) attacks over AES implementation. In [23], three attack scenarios are presented: single-bit DPA, multi-bit DPA, and a correlation analysis. The authors implement AES in two platforms. The study concludes by proposing an enhanced DPA approach that strategically organizes plaintext inputs to maximize the leakage. Additionally, authors in [24, 25] present AES key extraction attacks using power analysis on the ATmega328P microcontroller. Further, the authors in [26] present a power analysis attack specifically targeted against ASIC implementations of the AES. The authors in [27, 28] present their findings on power side-channel attacks targeting the advanced modes of AES. Additional works, such as [15, 14, 13], use the static power side channel by analyzing the static power consumption of a device. While impedance is a contributing factor in static power consumption, these studies focus exclusively on the effects of impedance changes as they relate to power consumption.
While impedance has always been a factor in SCA, it has only recently gathered attention in the realm of hardware security [29, 30, 12]. Prior to its recognition for cryptographic key recovery, many studies focused primarily on other factors. For example, the authors in [31] introduce a detection technique using board level impedance to detect physical modifications of the board. The study in [29] explores a method for non-destructively detecting hardware Trojans through internal impedance variations. The authors in [32], study a clustering method to classify ICs into Trojan-free and Trojan-infected groups. In addition to the usage of impedance in physical modification detection, it is recently attracting the focus of device switching activities. Authors in [19] use impedance measurements to disassemble software instruction types for anomaly monitoring and software integrity verification. The authors in [33] use the impedance of digital circuits to introduce a novel RFID tag design. In [11], the study of impedance changes induced by the switching activities is used to detect and identify the firmware. Authors in [12] use the impedance side channel to extract the AES key from the registers of an FPGA. These works present the run-time impedance variations of the devices as the side-channel.
However, to the best of our knowledge, no unified setup has been established to compare the sensitivities of impedance and power side channels. Therefore, this work primarily focuses on studying this sensitivity.
@@ III Power vs Impedance: Key Extraction
In this section, we discuss the attack methodology used against the AES implementation, focusing on the exploitation of both power and impedance side channels. We detail the threat model, followed by the leakage models used in the attack. Additionally, we discuss the metrics used to evaluate the effectiveness of both power and impedance side channels.
@@ III-A Source of Impedance Side-channel
Impedance SCA introduces a novel approach to probing the intricate and often overlooked behaviors of chips, particularly in memory chips. This technique methodically explores memory operations, revealing how variations in impedance could potentially disclose information about the data being processed and stored. Central to this analysis is the understanding that multiple physical and architectural factors influence the impedance characteristics of a chip [19, 29].These factors include parasitic elements inherent to the physical design, interconnection structures, variations in transistor properties across the silicon substrate, and the distinct semiconductor material characteristics associated with different transistor configurations.
In memory chips such as static random access memory (SRAM) and dynamic random access memory (DRAM), the control circuit manages memory access and data read/write operations. The memory array, organized in rows and columns, stores data at specific address locations. Three primary components define the impedance profile of memory: (i) the architecture and layout of the memory array, (ii) the cell type, and (iii) the associated data and control logic circuitry. Manufacturers tailor the arrangement of memory cells in rows and columns, as well as select the cell type (e.g., four-transistor (4T) or six-transistor (6T) cell structures in SRAM), based on factors such as performance requirements, power consumption, and area constraints, all of which influence the impedance characteristics.
In our study, we specifically focus on the impedance changes that correspond to the content stored in memory cells. The process of writing data to or reading from memory cells initiates a series of events that create unique impedance profiles for the bit sequences stored within the memory array. Factors such as transistor activation, variations in the fabrication process, and specific data retention logic used contribute to these distinctive impedance profiles. Furthermore, the profiles are further influenced by the nature of the data stored in the cells, their locations within the array, and the varying lengths of interconnections for each cell. Thus, impedance measurements are expected to reflect these variations, making them susceptible to information leakage.
@@ III-B1 Target Intermediate
The key extraction of the AES encryption using physical SCA usually targets the first or final round of the algorithm. The targeted round is used to model the leakage signal. In our study, we perform impedance SCA on the first round of the AES algorithm, targeting the output of the S-Box. The process begins with a bitwise XOR operation between the plaintext and the secret key, followed by the substitution step using the S-Box. The output value of the substitution step is stored in memory, and it is our targeted computational intermediate value. The steps of AES encryption are illustrated in Fig. 1.
In our hypothetical leakage model, we use the plaintext $p_{i}$, and the corresponding round subkey $k_{j}$ to define the targeted computational intermediate $m_{i}$, as presented in Eq. 3. This intermediate value is influenced by both the input plaintext and the secret cryptographic key, establishing a correlation between the side-channel leakages and the secret cryptographic key. Attackers can exploit the statistical properties of the leakage traces to extract the cryptographic key. By focusing on the first round and the S-Box output while effectively modeling the leakage signal, we aim to enhance the understanding of how impedance SCA can compromise cryptographic security.